OpenStudy (anonymous):

Use the product rule to find the derivative and the find the equation of the tangent line for the function f(x)=(x^2+x-3)(3x^3-8x+1 at the point x=-1

OpenStudy (anonymous):

The product rule is d(f(x)*g(y))=f(x)*g'(y)+g(y)*f'(x) Can you do the rest?

OpenStudy (anonymous):

do I just plug in the numbers

OpenStudy (anonymous):

Plugging x=-1 into the derivative will give you the corresponding y coordinate. You need to find enough information to be able to find an equation for the tangent line.

OpenStudy (anonymous):

so it would be \[(x^2+x-3)(3x^3-8x+1)=(x^2+x-3)(3x^3-8x+1)+(3x^3-x+1)(x^2+x-3) Right ?\]

OpenStudy (anonymous):

I'm not quite sure. Can you resend just the right hand side?

OpenStudy (anonymous):

\[(x^2+x-3)(3x^3-8x+1)=\]

OpenStudy (anonymous):

Yeah, so what's the derivative of that?

OpenStudy (anonymous):

I'm not sure what to do

OpenStudy (anonymous):

Let's say that \[f =x^2+x-3\] and \[g=3x^3-8x+1\] The product rule says the derivative of \[f \times g=f \times g \prime + g \times f \prime\] Let me know what you get